### v2

# Sharpness-Controlled Group Relative Policy Optimization with Token-Level Probability Shaping

Abstract
Reinforcement learning with verifiable rewards (RLVR) has become a practical route to improve large language model reasoning, and Group Relative Policy Optimization (GRPO) is a widely used optimizer in this setting. This paper revisits GRPO from a generalization perspective. Recent analysis shows that population performance can be controlled by a robust empirical objective that decomposes into the training loss plus a sharpness term measured by the gradient norm. We develop a token-level view of this sharpness term and show that GRPO can be dominated by a small subset of tokens with disproportionately large per-token gradients, which increases sharpness and can harm generalization. Motivated by this view, we propose Token-Regulated GRPO (TR-GRPO), which introduces a monotone probability shaping function to assign token weights based on the model’s own token probabilities, and integrates these weights into the standard GRPO. Our analysis yields a bound that isolates a probability dependent multiplicative factor in token-gradient magnitudes, explaining how probability-aware weighting suppresses sharp directions while preserving learning signal on semantically critical tokens. Experiments on logic puzzles, mathematical reasoning, and tool-augmented question answering show consistent improvements over GRPO, along with smoother gradient-norm trajectories, supporting TR-GRPO as a simple and effective generalization-oriented upgrade to GRPO for RLVR.
1 Introduction
Large language models (LLMs) have recently demonstrated remarkable progress in complex reasoning tasks such as mathematics and programming, with systems like OpenAI’s O-series (jaech2024openai), DeepSeek-R1 (guo2025deepseek), Kimi K2 (team2025kimi), and Qwen3 (yang2025qwen3) achieving state-of-the-art performance. A central driver of these advances is Reinforcement Learning with Verifiable Rewards (RLVR) (gao2024designing; lambert2024tulu; team2025kimi; guo2025deepseek; yang2025qwen3), which commonly employs reinforcement learning on an LLM with a rule-based outcome reward, such as a binary indicator reward for mathematical or logical validity. RLVR has emerged as a practical framework for strengthening the reasoning abilities of LLMs, offering stable and task-aligned reward signals without the need for costly human annotation or additional models. Among the RL algorithms used in this setting, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) has become a popular choice due to its simplicity and strong empirical results.
Although recent progress in RLVR has been driven by new algorithms (yu2025dapo; yue2025vapo; guan2025rstar), cross-domain applications (xue2025dancegrpo; liu2025flow; pan2025medvlm), and unexpected empirical observations  (wang2025reinforcement; yue2025vapo; zhao2025absolute), current implementations typically still treat all tokens equally without considering which ones are truly important, overlooking the fact that model responses often contain a mixture of both high-quality and low-quality tokens, especially as response length increases. Such uniform training disregards the varied functional roles tokens play in reasoning, which can hinder performance by neglecting critical tokens and amplifying noise from less relevant ones.
(a) Frequent tokens with the highest average probability.
(b) Frequent tokens with the lowest average probability.
Figure 1: Word clouds of the top 100 high- vs. low-probability tokens selected from frequently occurring words. High-probability tokens (left) primarily consist of mathematical and logical operators, brackets, and variable names, where even small errors can invalidate an entire solution, whereas low-probability tokens (right) mostly consist of generic content words that are less critical.
Motivated by these observations, we ask: which tokens actually drive the update during RLVR ?. Under the generalization view in Theorem 3.1 and the sharpness surrogate in Eq. (8), we aim to regulate token-wise updates to reduce sharp directions while preserving learning signal on semantically critical tokens; To check whether this aligns with semantic importance, we visualize the top 100 frequent high probability tokens and low-probability tokens ranked by their average model probability $\bar{\pi}(o_{t})=\tfrac{1}{\#\mathrm{occ}(o_{t})}\sum_{o_{t}}\pi_{\theta}(o_{t})$ (Figures 1(a)1(b)) where $\mathrm{occ}(o_{t})$ denotes the number of occurrences of token $o_{t}$. Figure 1(a) shows that high-probability tokens concentrate on mathematical and logical structure—operators, brackets, variable names, and formatting markers—where even a single error can break the entire solution, whereas low-probability tokens are mostly generic content words (e.g., “output,” “particular,” “location”) that contribute less to the core reasoning and are more easily replaced without changing meaning. This creates a clear tension: standard GRPO naturally magnifies low-probability tokens (via larger gradients) even though high-probability tokens carry the most critical signal for correctness. We therefore need a mechanism to mitigate over-updates on low-probability tokens while preserving and even reinforcing updates on high-probability ones.
To tackle this, we propose a simple yet effective method: Token-Regulated GRPO, which introduces a probability-aware token weight that increases with the token’s probability as described in Eq. (10), so that high-probability tokens are emphasized and low-probability ones are downweighted. Experiments across logic puzzles, mathematical reasoning, and agentic QA confirm that Token-Regulated GRPO consistently outperforms GRPO. In particular, across all these RLVR settings, TR-GRPO delivers substantial gains: on the K&K logic puzzles benchmark, Qwen2.5-3B improves from 0.39 to 0.63 (Table 1); in agentic multi-hop QA, exact-match accuracy nearly doubles from 13.84 to 27.29 (Table 3); and on math benchmarks including OlympiadBench and Minerva, we observe average improvements of +6–10% (Table 4.2.2). Beyond raw accuracy, TR-GRPO yields smoother gradient trajectories and more stable training dynamics (Figure 3), confirming the benefit of probability-aware reweighting.
2 Background
2.1 Large Language Models
Most Large Language Models (LLMs) adopt a transformer decoder-only architecture (vaswani2017attention), parameterized by $\theta\in\mathbb{R}^{d}$ and denoted as $\pi_{\theta}$. The basic unit of an LLM is the token, which may correspond to a word, subword, or character, and is selected from a finite vocabulary $\mathcal{V}=\{v^{1},\dots,v^{N}\}$ of size $N$. Given a prompt $q$, the model generates a sequence of tokens $o=(o_{1},\dots,o_{T})$ autoregressively. At each step $t$, the model produces a distribution over $\mathcal{V}$ conditioned on $q$ and the previously generated tokens $o_{<t}$, from which the next token is sampled:
$o_{t}\sim\pi_{\theta}(\cdot\mid q,o_{<t}).$ (1)
This process continues until an end-of-sequence (EOS) token is produced or the sequence length reaches a predefined maximum $t_{\max}$.
While pretrained and supervised fine-tuned LLMs achieve fluency, practical applications often require alignment with human preferences or reasoning ability, which is not easily captured by likelihood training alone. To address this, one may introduce a reward function $r(q,o)$ that quantifies the quality of a generated sequence $o$ given prompt $q$. The learning problem can then be cast as reinforcement learning, where the prompt $q$ represents the state, each generated token is an action, and the full output $o$ yields a terminal reward. Accordingly, the optimization objective is formulated as: $\max_{\theta}\;\mathbb{E}_{q\sim\mathcal{Q},\,o\sim\pi_{\theta}(\cdot\mid q)}\big[\,r(q,o)\,\big],$ where $\mathcal{Q}$ denotes a dataset of prompts.
2.2 Reinforcement Learning with Verifiable Rewards (RLVR)
Reinforcement Learning with Verifiable Rewards (RLVR)  (gao2024designing; lambert2024tulu; team2025kimi; guo2025deepseek; yang2025qwen3) provides a framework for optimizing LLMs in domains where correctness can be deterministically verified. Typical applications include code generation validated by unit tests, mathematical problem solving with symbolic checkers, and factual QA with exact-match rules or programmatic validators. Unlike preference-based reinforcement learning, RLVR leverages a predefined verifier to produce a reward without requiring human labels. Formally, given a dataset of prompts $\mathcal{D}$, a policy $\pi_{\theta}$, and a frozen reference model $\pi_{\mathrm{ref}}$, the objective is:
$\max_{\theta}\;\mathbb{E}_{q\sim\mathcal{Q},\,o\sim\pi_{\theta}(\cdot\mid q)}\Big[R_{\phi}(q,o)\Big]-\beta\,\mathbb{D}_{\mathrm{KL}}\!\left[\pi_{\theta}(o\mid q)\,\|\,\pi_{\mathrm{ref}}(o\mid q)\right],$ (2)
where $R_{\phi}$ is the verifiable reward function and $\beta$ controls the KL regularization. A typical verifiable reward function is defined as:
$R_{\phi}(q,o)=\begin{cases}1&\text{if }\mathrm{match}(o,o_{g}),\\[4.0pt] -1&\text{otherwise},\end{cases}$ (3)
where $o_{g}$ is the ground-truth answer and $\mathrm{match}(\cdot,\cdot)\in\{0,1\}$ indicates whether the generated output matches. More generally, $R_{\phi}$ may be graded (e.g., partial credit, length penalty, latency) while remaining verifiable by rule-based approaches (wang2024openhands) or model-based verifiers (ma2025general).
In summary, RLVR provides a principled setting where optimization is guided by verifiable correctness, making it a natural foundation for training reasoning-oriented LLMs.
2.3 Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO), first introduced by shao2024deepseekmath, is a policy-gradient algorithm widely used for optimizing the objective in Eq. 2. Unlike the popular PPO algorithm  (schulman2017proximal), which requires training an additional value function alongside the LLM, GRPO removes the dependency on an explicit value model. Instead, it estimates advantages by normalizing rewards within a group of sampled responses to the same prompt. Specifically, for a prompt $q$ with $G$ sampled responses $\{o_{i}\}_{i=1}^{G}$ and associated scalar rewards $\{r_{i}\}_{i=1}^{G}$, GRPO defines a group-normalized advantage as:
$\hat{A}_{i,t}=\frac{r_{i}-\mathrm{mean}(\{r_{j}\}_{j=1}^{G})}{\mathrm{std}(\{r_{j}\}_{j=1}^{G})},$ (4)
The effectiveness of the above normalization method can be interpreted through the lens of reward shaping: by emphasizing relative differences among candidate outputs for the same prompt, it strengthens the stability of the gradient signal and maintains its reliability, even under sparse reward conditions (hu2020learning). In this study, we employ a variant of GRPO to optimize the policy model $\pi_{\theta}$. The optimization objective of GRPO can be expressed as follows:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}\big[q\!\sim\!\mathcal{Q},\;\{o_{i}\}_{i=1}^{G}\!\sim\!\pi_{\theta_{\mathrm{old}}}(O\mid q)\big]$ (5)
$\quad\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left\{\min\!\big(r_{i,t}(\theta)\,\hat{A}_{i,t},\,\operatorname{clip}(r_{i,t}(\theta),\,1-\epsilon_{l},\,1+\epsilon_{h})\,\hat{A}_{i,t}\big)\;-\;\beta\,\mathbb{D}_{\mathrm{KL}}\!\left[\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\right]\right\},$
where $r_{i,t}(\theta)=\tfrac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(o_{i,t}\mid q,o_{i,<t})}$ is the importance sampling ratio and the KL divergence term is defined as:
$\mathbb{D}_{\mathrm{KL}}\!\left[\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\right]=\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}-\log\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}-1.$ (6)
Here, $\pi_{\theta_{\mathrm{old}}}$ denotes the policy used to sample responses, $\pi_{\mathrm{ref}}$ is a frozen reference model, and $\epsilon_{l},\epsilon_{h}$ are the PPO-style clipping threshold hyper-parameters and $\beta$ controls KL regularization.
3 Methodology
3.1 A generalization view of GRPO
GRPO-style RLVR trains a policy $\pi_{\theta}$ using samples $(q,o_{1:G})$ drawn from a joint distribution $p(q,o_{1:G})\triangleq\mathcal{Q}(q)\prod_{i=1}^{G}\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)$. In practice, optimization is performed on a finite dataset $\mathcal{S}=\{(q_{k},o_{k1:G})\}_{k=1}^{N}$, which creates a gap between the population/generalization loss $\mathcal{L}_{\mathcal{D}}(\pi_{\theta})$ (over $p$) and the empirical loss $\mathcal{L}_{\mathcal{S}}(\pi_{\theta})$ (over $\mathcal{S}$).
A key theoretical message from recent GRPO generalization analysis is that, with high probability over the sampled dataset, the population loss can be controlled by a robust version of the empirical loss:
Theorem 3.1 (Generalization via local robustness).
With probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$,
$\mathcal{L}_{\mathcal{D}}(\pi_{\theta})\;\leq\;\max_{\|\theta^{\prime}-\theta\|\leq\rho}\;\mathcal{L}_{\mathcal{S}}(\pi_{\theta^{\prime}})\;+\;\text{(lower-order terms)}.$ (7)
Moreover, a first-order expansion around $\theta$ yields the approximation
$\max_{\|\theta^{\prime}-\theta\|\leq\rho}\mathcal{L}_{\mathcal{S}}(\pi_{\theta^{\prime}})\;\approx\;\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\;+\;\rho\,\|\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\|_{2},$ (8)
which highlights two complementary targets for better generalization: (i) reducing the empirical loss, and (ii) reducing the sharpness term measured by the gradient norm. This section develops a token-level view of how GRPO shapes $\|\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\|_{2}$, and proposes a probability-aware reweighting that directly suppresses sharp directions while preserving learning signal on semantically critical tokens.
3.2 Token-level gradients in GRPO
We start from the derivative of the GRPO objective with respect to model parameters:
$\nabla_{\theta}J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}}$
$\Biggl[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\underbrace{\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\hat{A}_{i,t}\cdot\mathbb{I}_{\text{trust}}\!\left(\tfrac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)+\beta\tfrac{\pi_{\text{ref}}(o_{i,t})}{\pi_{\theta}(o_{i,t})}-\beta\right)}_{\gamma_{i,t}}\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\Biggr].$ (9)
For simplicity, $\pi_{\theta}(o_{i,t}):=\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})$. The trust-region indicator is
$\mathbb{I}_{\text{trust}}\!\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)=\begin{cases}0,&\hat{A}_{i,t}>0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>1+\epsilon_{h},\\[4.0pt] 0,&\hat{A}_{i,t}<0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<1-\epsilon_{l},\\[4.0pt] 1,&\text{otherwise}.\end{cases}$
A crucial observation is that the sharpness term in Eq. (8) depends on the magnitude of token gradients. Under standard smoothness assumptions on transformer layers, one can bound $\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}$ and show it scales with $(1-\pi_{\theta}(o_{i,t}))$: low-probability tokens induce larger gradients, hence contribute more to sharpness, even when such tokens are often semantically replaceable in natural language reasoning traces.
3.3 Token-Regulated GRPO for better generalization
Motivated by Eq. (7)-(8), we aim to reduce sharpness by explicitly controlling token-level gradient magnitudes. Following the generalization development, we introduce an increasing probability shaping function $\omega(\cdot)$ and define a token weight $w_{i,t}:=\omega(\pi_{\theta}(o_{i,t}))$. This yields a weighted GRPO surrogate that remains compatible with PPO/GRPO clipping and advantage learning, while steering optimization toward flatter (less sharp) regions.
We propose Token-Regulated GRPO (TR-GRPO):
$\mathcal{J}_{\text{TR-GRPO}}(\theta)$ $=\mathbb{E}\!\big[q\!\sim\!\mathcal{Q},\;\{o_{i}\}_{i=1}^{G}\!\sim\!\pi_{\theta_{\mathrm{old}}}(O\mid q)\big]$ (10)
$\quad\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\Bigl\{\min\!\bigl(w_{i,t}r_{i,t}(\theta)\hat{A}_{i,t},\,\operatorname{clip}(w_{i,t}r_{i,t}(\theta),1-\epsilon_{l},1+\epsilon_{h})\hat{A}_{i,t}\bigr)$
$\quad{}\hskip 80.00012pt-\;\beta\,\mathbb{D}_{\mathrm{KL}}\!\bigl[w_{i,t};\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\bigr]\Bigr\},$
where $r_{i,t}(\theta)\triangleq\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(o_{i,t}\mid q,o_{i,<t})}$ and
$\mathbb{D}_{\mathrm{KL}}\!\left[w_{i,t};\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\right]=w_{i,t}\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}-\log\!\left(w_{i,t}\frac{\pi_{\mathrm{ref}}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}\right)-1.$ (11)
We instantiate $w_{i,t}$ with a monotone, probability-aware mapping and stop-gradient to avoid the model gaming the weights:
$w_{i,t}\;=\;\operatorname{clip}\!\Big(\alpha\cdot\big[\sigma\!\big(\tfrac{\text{sg}[\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})]}{\tau}\big)-\mu\big],\;L,\;U\Big),$ (12)
where $\text{sg}[\cdot]$ is the stop-gradient operator.
The token-weighted surrogate preserves the GRPO gradient structure up to the multiplicative factor $w_{i,t}$. Concretely, the token-wise contribution to $\nabla_{\theta}\mathcal{J}_{\text{TR-GRPO}}$ can be written as
$g_{i,t}\;=\;\gamma_{i,t}\,w_{i,t}\,\nabla_{\theta}\log\pi_{\theta}(o_{i,t}),$ (13)
with a modified trust-region indicator:
$\mathbb{I}_{\text{trust}}\!\left(\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})},\hat{A}_{i,t}\right)=\begin{cases}0,&\hat{A}_{i,t}>0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>w_{i,t}^{-1}(1+\epsilon_{h}),\\[4.0pt] 0,&\hat{A}_{i,t}<0\ \text{and}\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<w_{i,t}^{-1}(1-\epsilon_{l}),\\[4.0pt] 1,&\text{otherwise}.\end{cases}$
Under the same layer-wise smoothness and bounded Jacobian assumptions used in standard transformer analyses, we can bound $\|g_{i,t}\|_{2}$ and isolate the key multiplicative term $w_{i,t}(1-\pi_{\theta}(o_{i,t}))$:
Theorem 3.2 (Token gradient bound for TR-GRPO).
Under the same assumptions as the standard bound on $\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}$, the token-gradient norm satisfies
$\frac{w_{i,t}(1-\pi_{\theta}(o_{i,t}))}{\sqrt{L}}\,|\gamma_{i,t}|\sum_{l=1}^{L}\left(a^{W}a_{l}^{G}\prod_{j=l}^{L}a_{j}^{J}\right)\leq\|g_{i,t}\|_{2}\leq\sqrt{2}\,w_{i,t}(1-\pi_{\theta}(o_{i,t}))\,|\gamma_{i,t}|\sum_{l=1}^{L}\left(b^{W}b_{l}^{G}\prod_{j=l}^{L}b_{j}^{J}\right).$ (14)
For low-probability tokens, $(1-\pi_{\theta}(o_{i,t}))$ is large while $w_{i,t}$ is small; for high-probability tokens, $(1-\pi_{\theta}(o_{i,t}))$ is small while $w_{i,t}$ is large. Thus, TR-GRPO stabilizes the product $w_{i,t}(1-\pi_{\theta}(o_{i,t}))$ across tokens, suppressing extreme token gradients that would otherwise dominate $\|\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(\pi_{\theta})\|_{2}$. By Eq. (8), reducing this gradient norm directly reduces sharpness, which tightens the robust empirical bound in Eq. (7) and improves generalization.
To understand what high- vs. low-probability tokens look like in practice, we visualize the top 100 tokens by average model probability. High-probability tokens are dominated by mathematical and logical operators, brackets, variable names, and formatting markers; even small errors in these tokens can invalidate an entire solution. Low-probability tokens are mostly generic content words that contribute less to the core reasoning and are more easily replaceable. This observation motivates emphasizing high-probability tokens, since they carry the most critical signal for correctness.
To determine whether emphasizing high-probability tokens is indeed beneficial, we compare: (i) TR-GRPO, which assigns higher weights to high-probability tokens and lower weights to low-probability ones; and (ii) a Reverse variant that inverts TR-GRPO’s weight by assigning 2w2-w to any token whose TR-GRPO weight is $w$. As shown in Figure 2, TR-GRPO (blue) consistently outperforms both Reverse (orange) and the GRPO baseline (green) across K&K puzzle sizes (3–7), while Reverse tracks GRPO closely without clear improvement. Beyond accuracy, this ablation supports the generalization view: prioritizing semantically critical, high-probability tokens improves robustness of the learned behavior, whereas over-amplifying rare tokens does not.
Figure 2: Accuracy on the K&K Logic Puzzles benchmark, broken down by puzzle size (3–7 people). TR-GRPO consistently achieves higher accuracy than GRPO across all difficulty levels, while the Reverse variant that emphasizes low-probability tokens yields performance comparable to GRPO without clear improvement.
Eq. (8) suggests that smaller gradient norms correspond to less sharp updates and hence better generalization. From Theorem 3.2, TR-GRPO suppresses extreme token gradients through the stabilizing factor $w_{i,t}(1-\pi_{\theta}(o_{i,t}))$. To validate this mechanism, we track gradient norms during training under three RLVR settings in Figure 3. Compared to GRPO, TR-GRPO exhibits smaller fluctuations and fewer outlier spikes, consistent with reduced sharpness and improved generalization.
Figure 3: Gradient norm trajectories during training under GRPO vs. TR-GRPO across three RLVR settings. TR-GRPO consistently exhibits lower variability and fewer spikes than GRPO, consistent with reduced sharpness in Eq. (8) and the bound in Eq. (14).
In addition to gradient-norm analysis, we also track training rewards across logic, math, and agentic settings. TR-GRPO yields higher and more stable reward trajectories than GRPO, indicating more effective learning while simultaneously controlling sharpness.
Figure 4: Training reward trajectories during training under GRPO vs. TR-GRPO across three RLVR settings. TR-GRPO achieves higher reward while also exhibiting lower sharpness as reflected by gradient norms.
4 Experiments
We conduct extensive experiments across multiple RLVR benchmarks to assess the effectiveness of TR-GRPO. The results demonstrate that our method consistently outperforms GRPO, delivering stronger reasoning ability and more stable training dynamics.
4.1 Experimental Setup
To validate the effectiveness and generality of our proposed method, we conduct experiments in three widely used RLVR settings that stress different aspects of reasoning: (i) Logic, (ii) Math, and (iii) Agentic.
For the logic data, we adopt the K&K (Knights and Knaves) logic puzzles introduced in (xie2024memorization). Each puzzle describes a set of people who always lie or always tell the truth, and the task is to determine their identities given a set of statements. For training, we use the same dataset splits as in (yang2025not), which include synthetic instances of increasing difficulty from 3 to 7 (measured by the number of people per puzzle). The reward function checks both the output format (requiring explicit <think> and <answer> tags) and the correctness of the final assignment. During evaluation, we follow the official harness and report accuracy across all difficulty levels, as well as breakdowns by puzzle size. This setting emphasizes step-by-step deductive reasoning and tests whether the model can align reasoning traces with verifiable logical consistency.
We next consider mathematical reasoning tasks, where correctness can be verified automatically against gold-standard answers. Following  (yang2025not), we train on datasets that combine symbolic manipulation and arithmetic word problems, using a binary reward signal: $1$ if the final boxed answer matches the reference solution and 0 otherwise. The training corpus includes diverse math reasoning prompts designed to encourage structured derivations rather than direct guessing. For evaluation, we adopt multiple benchmarks that are standard in math-focused RLVR research: Olympiad Bench  hendrycks2021measuring, Minerva  (lewkowycz2022solving), MATH-500  (he2024olympiadbench), AMC 2022-2023 and AIME 2024. For the first three benchmarks, evaluation is conducted using greedy decoding. For the last two benchmarks, consistent with standard practice, we generate 16 responses per question and report the mean accuracy across these samples (avg@16). Notably, since AIME 2024 is an extremely challenging dataset, we also report pass@16, which counts a problem as solved if at least one of the 16 responses is correct.
For agentic reasoning tasks, we examine an agentic setting that requires knowledge-intensive reasoning augmented with retrieval tools. Following the Search-R1  (jin2025search) configuration from VerlTool framework  (jiang2025verltool), the model interacts with an external tool server providing a retriever and other utilities such as Python or SQL. Training data consists of open-domain QA datasets where each instance requires grounding reasoning steps with retrieved evidence. The reward is based on exact-match correctness of the final answer, while intermediate tool calls are masked in the policy loss to avoid leakage of supervision. For evaluation, we follow the VerlTool benchmark and report exact-match accuracy on both General Q&A benchmarks (NQ  (kwiatkowski2019natural), TriviaQA  (joshi2017triviaqa), PopQA (mallen2022not)) and multi-hop Q&A benchmark (HotpotQA (yang2018hotpotqa), 2Wiki (ho2020constructing), MuSiQue (trivedi2022musique), Bamboogle (press2022measuring)).
We leave further implementation details and hyperparameters to Appendix B with detailed evaluation setting adopted by our experiments.
4.2 Main Results
Besides the detailed analysis of low and high-probability tokens in Section 3, we now present the main empirical results of our study. Across all three widely used RLVR settings include logic, math, and agentic, our proposed TR-GRPO method achieves substantial and consistent gains over the widely used GRPO baseline.
4.2.1 Experiments on K&K Logic Puzzles
Table 1 reports results on the K&K Logic Puzzles benchmark, where TR-GRPO achieves significant improvements over GRPO across all difficulty levels. Notably, TR-GRPO boosts the average accuracy of Qwen2.5-3B from 39% (with GRPO) to 63%, representing a relative gain of over 60%. Similarly, for Qwen2.5-7B, TR-GRPO increases the average accuracy from 77% to 91% (+18%). It is worth noting that this task remains challenging even for powerful proprietary models such as GPT-4o, DeepSeek-R1, and o1, which achieve average accuracies of only 36–78%.
Table 1: Experimental results on the K&K Logic Puzzles benchmark. The best results are indicated in bold.
Difficulty by Number of People
Model 3 4 5 6 7 Avg.
GPT-4o 0.57 0.49 0.32 0.23 0.21 0.36
o1-2024-12-17 0.51 0.38 0.38 0.35 0.30 0.38
Deepseek-R1 0.73 0.77 0.78 0.75 0.88 0.78
[…]
   + GRPO 0.64 0.47 0.35 0.30 0.21 0.39
   + TR-GRPO 0.76 0.76 0.61 0.59 0.44 0.63 $\uparrow$ 61.5%
[…]
   + GRPO 0.91 0.91 0.77 0.65 0.61 0.77
   + TR-GRPO 0.95 0.95 0.92 0.87 0.84 0.91 $\uparrow$ 18.2%
4.2.2 Experiments on Math-related Datasets
To further demonstrate the effectiveness of our method, we conducted experiments on Math-related datasets. Table 4.2.2 summarizes performance on mathematical reasoning datasets. TR-GRPO consistently surpasses GRPO, with average gains of 6–10% across OlympiadBench, Minerva, MATH-500. AMC, and AIME. In particular, on AIME, TR-GRPO raises pass@16 from 16.5% to 18.1% and average@16 from 28.9% to 43.3%, demonstrating stronger robustness in symbolic and arithmetic reasoning. These results clearly highlight the superiority of our proposed TR-GRPO over GRPO baseline in advancing mathematical reasoning.
Table 2: Experimental results on math-related datasets (DSR for DeepScaleR and ORZ for Open Reasoner-Zero). The best results are indicated in bold.
Dataset Algorithms Olympiad Bench Minerva MATH 500 AMC avg@16 AIME pass@16 AIME avg@16 Avg.
[…]
DSR + GRPO 36.50 29.66 74.67 47.72 28.89 16.46 38.98
+ TR-GRPO 38.48 32.35 79.40 46.84 43.33 18.13 43.09 $\uparrow$ 10.5%
ORZ + GRPO 38.23 27.69 78.33 49.57 32.22 12.92 39.83
+ TR-GRPO 40.12 30.51 78.60 45.48 43.33 16.88 42.49 $\uparrow$ 6.7%
4.2.3 Experiments on Agentic VerlTool Search
To extend our evaluation, we further examine TR-GRPO on agentic RLVR tasks. Table 3 presents results in the agentic RLVR setting, where TR-GRPO delivers the largest improvements. On knowledge-intensive QA tasks, TR-GRPO nearly doubles the performance of GRPO, improving average exact-match accuracy from 13.8% to 27.3% (+97%) for Qwen2.5-3B, and yielding a 14% relative gain on Qwen3-4B-Instruct-2507. These results confirm the effectiveness of token-level regulation in tool-augmented reasoning, where stability and precision are critical. Together, these results establish TR-GRPO as a consistently stronger optimization method than GRPO across diverse RLVR domains, with particularly striking gains in logic puzzles and agentic reasoning.
Table 3: Experimental results for the agentic task using VT-Search on knowledge-QA benchmarks. $\dagger$ represents in-domain datasets and $\star$ represents out-domain datasets. The best results are indicated in bold.
Model General QA Multi-hop QA
NQ TriviaQA PopQA HQA 2Wiki Musique Bamboogle Avg.
[…]
+ GRPO 13.50 29.40 11.50 14.50 16.30 1.30 10.40 13.84
+ TR-GRPO 41.80 52.30 39.10 24.50 20.30 5.00 8.00 27.29 $\uparrow$ 97.2%
[…]
+ GRPO 46.48 59.84 40.39 35.54 28.87 10.26 27.20 35.21
+ TR-GRPO 48.23 64.11 46.00 36.60 30.89 10.26 41.60 40.18 $\uparrow$ 14.1%
4.3 Ablation studies
To better understand the role of token-level weighting in our framework, we conduct an ablation study on the K&K Logic Puzzles benchmark using Qwen2.5-3B-Instruct.
Table 4: Ablation study for token weight estimation on Qwen2.5-3B-Instruct on the K&K Logic Puzzles benchmark. The best results are indicated in bold.
Difficulty by Num of People Avg.
Model 3 4 5 6 7
[…]
Equal weight 0.64 0.47 0.35 0.30 0.21 0.39
Random weight 0.67 0.43 0.37 0.32 0.24 0.41
Reverse weight 0.59 0.46 0.35 0.33 0.20 0.39
[…]
We compare TR-GRPO against several alternative weighting strategies, including: (i) Equal weight, which assigns every token a constant value of $1$; (ii) Random weight, where each token weight is sampled uniformly from $[0.5,1.5]$; and (iii) Reverse weight, which inverts the weights computed by our method. Specifically, if TR-GRPO assigns a weight $w$ to a token $t$ in sequence $y$, the reverse weighting assigns it 2w2-w. As shown in Table 4, TR-GRPO consistently achieves the best performance across all difficulty levels. Equal, Random, and Reverse weighting provide only modest gains, but are clearly outperformed by our probability-aware design.
5 Conclusion
In this work, we introduced TR-GRPO, a simple yet effective extension of GRPO that incorporates token-level weighting to enhance stability and effectiveness in reinforcement learning for reasoning. By assigning probability-aware token weights, TR-GRPO stabilizes gradient updates and ensures that critical high-probability tokens are emphasized. Our theoretical analysis established how token probabilities govern gradient norms, and our extensive experiments across logic, math, and agentic reasoning confirmed substantial and consistent improvements over the GRPO baseline. Together, these findings highlight the importance of regulating token-level contributions in reinforcement learning, and position TR-GRPO as a robust framework for enhancing the stability and reasoning capability of large language models.
Limitations.
A primary limitation of TR-GRPO lies in the additional computational overhead introduced by token-level weighting. Specifically, each update requires estimating and applying weights for all generated tokens, which slightly increases the per-step training cost compared to standard GRPO. Nevertheless, as shown in Appendix C.2, the overhead remains within a practically acceptable range and does not hinder scalability to larger models or datasets.
Appendix A Related Work
Large-Scale Reasoning Models.
Large language models (LLMs)  (lambert2024tulu; gao2024designing; team2025kimi; guo2025deepseek; yang2025qwen3) have recently achieved remarkable progress across a variety of natural language processing tasks. In particular, researchers are increasingly focused on enhancing their performance in reasoning-heavy domains such as mathematics (cobbe2021training; hendrycks2024measuring), coding (jain2024livecodebench), and scientific reasoning (rein2024gpqa). rStar-Math (guan2025rstar) develops a self-evolved deep-thinking strategy that significantly boosts the reasoning capabilities of smaller LLMs. Snell et al. (snell2024scaling) introduce dense, process-based verifier reward models that adaptively refine a model’s response distribution based on the test-time prompt to strengthen reasoning skills. OpenAI’s O-series (jaech2024openai) applies large-scale reinforcement learning to train models capable of solving complex reasoning tasks, achieving state-of-the-art results on multiple benchmarks.
Reinforcement Learning for Large Language Model.
Before the emergence of reasoning-focused systems such as OpenAI’s O-series (jaech2024openai), reinforcement learning (RL) had been widely applied in the form of reinforcement learning from human feedback (RLHF) to enhance large language models’ (LLMs) instruction-following and preference alignment (ouyang2022training). RLHF approaches are typically divided into online and offline optimization. Online algorithms—including PPO (schulman2017proximal), GRPO (shao2024deepseekmath), and REINFORCE (williams1992simple)—update the model by generating outputs during training and receiving immediate reward signals. Offline variants such as DPO (rafailov2023direct), SimPO (meng2024simpo), and KTO (ethayarajh2024kto) optimize policies using pre-collected preferences from annotators or LLMs. While offline methods are generally more efficient, they often trail online training in terms of performance (tang2024understanding). More recently, reinforcement learning with verifiable rewards (RLVR) has been introduced as a promising paradigm for improving LLM reasoning, especially in domains like mathematics and programming. OpenAI o1 (jaech2024openai) was the first to demonstrate that RL can successfully scale reasoning abilities. Subsequent models—including DeepSeek-R1 (guo2025deepseek), Kimi-2 (team2025kimi), and Qwen3 (yang2025qwen3)—have either matched or surpassed its performance. Notably, DeepSeek-R1 highlights how strong reasoning can emerge from outcome-based optimization using online RL, particularly GRPO (shao2024deepseekmath). These advances also inspired the “zero RL” paradigm, where reasoning capabilities are elicited directly from the base model without explicit RL fine-tuning. Building on this idea, follow-up work has proposed methods such as DAPO (yu2025dapo), VAPO (yue2025vapo), SimpleRLZoo (zeng2025simplerl), and Open-Reasoner-Zero (hu2025open), which further explore RL-based reasoning at scale.
Appendix B Implementation Details
B.1 Experiments setup
This appendix provides complete details of datasets, prompts, reward design, rollout/training configurations, and evaluation protocols for all three RLVR settings used in this paper: Logic (K&K), Math, and Agentic (VT-Search).
B.1.1 Logic: Knights & Knaves (K&K)
Following  xie2025logic; yang2025not, we adopt LLMs after instruction fine-tuning (Qwen2.5-3B-Instruct  (yang2025qwen3) and Qwen2.5-7B-Instruct-1M  (yang2025qwen2)) as the initialization point. The tailored prompt designed for the LLMs is provided below.
[…]
To promote chain-of-thought (CoT) reasoning in LLMs,  xie2025logic introduces a reward function with two main components, as shown in Table 5. The output is judged as fully correct if the LLMs generate CoT reasoning wrapped inside <think></think> tags, and the final prediction enclosed within <answer></answer> tags.
Table 5: Reward design for K&K (Logic-RL).
[…]
Completely Correct 1 2
Partially Correct 1-1 1.5-1.5
Completely Wrong 1-1 2-2
B.1.2 Math-related dataset
As mentioned in Section 4.1, inspired by  (yang2025not), we carry out additional experiments on two math-focused datasets, DSR-Uniform ($10,000$ problems, evenly covering difficulty levels) and ORZ ($57.000$ problems). In line with prior studies, we adopt Qwen2.5-7B  (yang2025qwen3) as the base model. In this setting, no instruction-tuned templates are applied; instead, we employ a simple prompt directly.
[…]
LLMs without post-training generally struggle to follow strict output formats. Consequently, format-related signals are not included during training. Moreover, math tasks usually admit only a single correct solution, making partial credit unnecessary. Hence, a binary reward scheme is sufficient: the model receives a reward of $1$ for a correct answer and 0 otherwise.
B.1.3 Agentic: VT-Search (knowledge-augmented QA)
Question answering tasks frequently demand access to external knowledge that goes beyond a model’s parametric memory, especially for factual queries and multi-hop reasoning. To address this, we follow the Search-R1  (jin2025search) configuration from VerlTool framework  (jiang2025verltool) which incorporates a FAISS-based retrieval module, allowing agents to query a local knowledge base and extract the most relevant evidence for answering complex questions.
Building on prior work (jin2025search; song2025r1), an E5 retriever (wang2022text) was employed with the 2018 Wikipedia dump (karpukhin2020dense) as the indexed corpus. The agent alternates between retrieval operations and reasoning steps to form complete answers. we adopt Qwen2.5-3B  (yang2025qwen3) and Qwen3-4B-Instruct-2507  (yang2025qwen3) as the base models.
For this task, we use accuracy as the main reward, defined as:
$R_{\text{search}}(\mathbf{x},\mathbf{y})=\begin{cases}1&\text{if }\text{match}(\mathbf{y},\mathbf{y}_{g})\\ -1&\text{otherwise}\end{cases}$ (15)
For evaluation, we follow the VerlTool benchmark and report exact-match scores on both General Q&A benchmarks (NQ  (kwiatkowski2019natural), TriviaQA  (joshi2017triviaqa), PopQA (mallen2022not)) and multi-hop Q&A benchmark (HotpotQA (yang2018hotpotqa), 2Wiki (ho2020constructing), MuSiQue (trivedi2022musique), Bamboogle (press2022measuring))
B.2 Hyperparameters
The key hyperparameter configurations for GRPO training are summarized in Table 6. We adopt the clip-higher technique from DAPO (yu2025dapo) to stabilize entropy and prevent collapse. Checkpoints are saved every 20 RL steps, and full implementation details are available in our code release. For token importance estimation (Eq. 12 in the main paper), we set $\alpha=2.0$ and $\mu=0.25$ with clipping bounds $L=1.0$ and $U=1.4$. The scaling parameter is fixed at $\tau=9.0$. All experiments are conducted on 8 × NVIDIA H100 GPUs.
Table 6: Key hyperparameters for GRPO training.
[…]
Group size per prompt $G$ 8 8 8
Sampling temperature 0.7 1.0 0.8
Max response length 4096 4096 4096
Optimizer / LR AdamW / $1\times 10^{-6}$ AdamW / $1\times 10^{-6}$ AdamW / $1\times 10^{-6}$
KL penalty coefficient 0.001 0.001 0.001
PPO clip ratio (low / high) 0.20   /   0.24 0.20   /   0.24 0.20   /   0.24
Mini-batch size 64 128 32
Micro-batch size (updates) 256 512 256
Appendix C Additional Experiment Results
C.1 Hyperparameter sensitivity analysis
In Eq. 12, the token weight is determined by five hyperparameters: $\alpha$, $\mu$, $L$, $U$, and $\tau$. Among these, $\alpha$ and $\mu$ mainly control the global mean of the weight distribution. Since the sigmoid term $\sigma(\text{sg}[\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})]/\tau)$ lies in $(0.5,1)$, we fix $\alpha=2$ and $\mu=0.25$, resulting in a centered range $(0.5,1.5)$ with mean value close to $1.0$. This ensures that token weights remain stable around unity, while still allowing enough variation to emphasize or de-emphasize tokens depending on their probability.
We then validate the two remaining hyperparameters, $\tau$ and the clipping range $(L,U)$. Table 7 reports results on the K&K Logic Puzzles benchmark using Qwen2.5-3B-Instruct. For $\tau$, we sweep across [0.5,2.0,7.0,9.0,10.0,20.0]{[0.5,2.0,7.0,9.0,10.0,20.0]}. The results indicate that performance is quite sensitive to $\tau$, with very small or very large values leading to degradation. The best accuracy is achieved in a moderate range around $\tau=9.0$, suggesting that excessive sharpening or flattening of the sigmoid output is suboptimal.
For $(L,U)$, we consider four ranges: $(1.0,1.2)$, $(1.0,1.4)$, $(0.8,1.5)$, and $(0.5,1.5)$. In contrast to $\tau$, the choice of clipping bounds has only a mild effect, and all settings achieve comparable performance. This suggests that the precise clipping range mainly acts as a safeguard against extreme values rather than a critical tuning factor.
Table 7: Hyperparameter sensitivity of token weight estimation on Qwen2.5-3B-Instruct on the K&K Logic Puzzles benchmark.
Parameter Value 3 4 5 6 7 Avg.
$\tau$ 0.5 0.63 0.48 0.35 0.31 0.20 0.39
2.0 0.64 0.49 0.37 0.33 0.19 0.40
7.0 0.67 0.62 0.41 0.37 0.24 0.46
9.0 0.76 0.76 0.61 0.59 0.44 0.63
10.0 0.77 0.79 0.58 0.57 0.37 0.62
20.0 0.74 0.68 0.51 0.46 0.35 0.55
($L,U$) (1.0, 1.2) 0.77 0.76 0.60 0.58 0.40 0.62
(1.0, 1.4) 0.76 0.76 0.61 0.59 0.44 0.63
(0.8, 1.5) 0.74 0.75 0.57 0.62 0.42 0.62
(0.5, 1.5) 0.70 0.65 0.59 0.60 0.44 0.60
C.2 Computational Costs
Table 8 reports the computational overhead of applying TR-GRPO compared to the GRPO baseline on the K&K Logic Puzzles dataset. The only additional operation required by TR-GRPO is the computation of token-level weights. Importantly, these weights are derived directly from the model’s own output probabilities and therefore do not require any auxiliary teacher models or extra forward passes. As a result, peak GPU memory usage remains essentially unchanged between GRPO and TR-GRPO across both the Qwen2.5-3B-Instruct and LLaMa3.1-8B-Instruct backbones.
Table 8: Computational cost comparison of GRPO and TR-GRPO on K&K Logic Puzzle Dataset.
Procedure Qwen2.5-3B-Instruct LLaMa3.1-8B-Instruct
GRPO TR-GRPO GRPO TR-GRPO
Training Time/Sample (minutes) 271.4 286.0 678.5 722.2
Peak GPU Memory (GB) 459.7 460.2 581.2 580.6
In terms of runtime, TR-GRPO does introduce a modest increase in training time per sample (e.g., 271.4 vs. 286.0 minutes on Qwen2.5-3B-Instruct, and 678.5 vs. 722.2 minutes on LLaMa3.1-8B-Instruct). However, this overhead is relatively minor compared to the substantial performance improvements reported in Section 4. Together, these results confirm that TR-GRPO delivers consistent gains in reasoning performance without imposing significant additional computational costs.
C.3 Cross-Model Validation
Beyond the Qwen family used in previous experiments—which has been the dominant backbone in recent reasoning research—we further validate TR-GRPO on other model families, including Mistral and LLaMA. In addition, we test on the base variant Qwen2.5-3B-Base instead of the usual instruction-tuned model. All experiments are conducted under the logic K&K setting, with results reported in Table 9.
Table 9: Experimental results on the K&K Logic Puzzles benchmark on various models. The best results are indicated in bold.
Difficulty by Number of People
Model 3 4 5 6 7 Avg.
[…]
   + GRPO 0.64 0.47 0.35 0.30 0.21 0.39
   + TR-GRPO 0.76 0.76 0.61 0.59 0.44 0.63 $\uparrow$ 61.5%
[…]
   + GRPO 0.91 0.91 0.77 0.65 0.61 0.77
   + TR-GRPO 0.95 0.95 0.92 0.87 0.84 0.91 $\uparrow$ 18.2%
[…]
   + GRPO 0.60 0.54 0.43 0.38 0.28 0.45
   + TR-GRPO 0.68 0.62 0.44 0.47 0.26 0.49 $\uparrow$ 8.9%
[…]
   + GRPO 0.29 0.16 0.09 0.11 0.03 0.14
   + TR-GRPO 0.47 0.27 0.18 0.15 0.08 0.23 $\uparrow$ 64.3%
[…]
   + GRPO 0.92 0.92 0.83 0.79 0.80 0.85
   + TR-GRPO 0.89 0.95 0.88 0.87 0.81 0.88 $\uparrow$ 3.5%
As shown in Table 9, across all models, TR-GRPO consistently outperforms GRPO, demonstrating that our method generalizes beyond Qwen-Instruct models. This provides further evidence that the proposed token-weighted strategy is broadly applicable and not limited to a single model family.
Appendix D Proof of Our Theories
Derivation of Derivative of GRPO in Eq. (3.2)
For $\hat{A}_{i,t}>0$, the inner term of the sum relevant to $o_{i,t}$ reduces to
$h\left(o_{i,t}\right)=\begin{cases}\left(1+\epsilon_{h}\right)\hat{A}_{i,t}&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>1+\epsilon_{h}\\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\hat{A}_{i,t}&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\leq 1+\epsilon_{h}\end{cases}$
The derivative w.r.t. $\theta$ becomes
$\nabla_{\theta}h\left(o_{i,t}\right)=\begin{cases}0&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}>1+\epsilon_{h}\\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\hat{A}_{i,t}&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\leq 1+\epsilon_{h}\end{cases}$
Combining with the KL term derivative, we gain
$\nabla_{\theta}h\left(o_{i,t}\right)+\beta\pi_{ref}\left(o_{i,t}\right)\frac{\nabla_{\theta}\pi_{\theta}\left(o_{i,t}\right)}{\pi_{\theta}\left(o_{i,t}\right)^{2}}-\beta\nabla_{\theta}\log\pi_{\theta}(o_{i,t})$
$=$ $\nabla_{\theta}h\left(o_{i,t}\right)+\beta\frac{\pi_{ref}\left(o_{i,t}\right)}{\pi_{\theta}\left(o_{i,t}\right)}\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)-\beta\nabla_{\theta}\log\pi_{\theta}(o_{i,t})$
For $\hat{A}_{i,t}<0$, the inner term of the sum relevant to $o_{i,t}$ reduces to
$h\left(o_{i,t}\right)=\begin{cases}\left(1-\epsilon_{l}\right)\hat{A}_{i,t}&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<1-\epsilon_{l}\\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\hat{A}_{i,t}&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\geq 1-\epsilon_{l}\end{cases}$
The derivative w.r.t. $\theta$ becomes
$\nabla_{\theta}h\left(o_{i,t}\right)=\begin{cases}0&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}<1-\epsilon_{l}\\ \frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\hat{A}_{i,t}&\frac{\pi_{\theta}(o_{i,t})}{\pi_{\text{old}}(o_{i,t})}\geq 1-\epsilon_{l}\end{cases}$
Combining with the KL term derivative, we gain
$\nabla_{\theta}h\left(o_{i,t}\right)+\beta\pi_{ref}\left(o_{i,t}\right)\frac{\nabla_{\theta}\pi_{\theta}\left(o_{i,t}\right)}{\pi_{\theta}\left(o_{i,t}\right)^{2}}-\beta\nabla_{\theta}\log\pi_{\theta}(o_{i,t})$
$=$ $\nabla_{\theta}h\left(o_{i,t}\right)+\beta\frac{\pi_{ref}\left(o_{i,t}\right)}{\pi_{\theta}\left(o_{i,t}\right)}\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)-\beta\nabla_{\theta}\log\pi_{\theta}(o_{i,t})$
Finally, leveraging the two above cases, we gain the final formula.
Lemma D.1.
Assume that $A_{1:m}$ is a sequence of matrices with $\sigma_{\text{min}}(A_{i})\geq a_{i}^{2}\geq 0$ and $\sigma_{\text{max}}(A_{i})\leq b_{i}^{2}$ for all $i\in\{1,...,m\}$. We then have
$\|x\|_{2}\prod_{i=m}^{1}a_{i}\leq\|x\prod_{i=m}^{1}A_{i}\|_{2}\leq\|x\|_{2}\prod_{i=m}^{1}b_{i}$
Proof.
We start with
$\|xA\|_{2}^{2}=xAA^{T}x^{T}.$
According to the Rayleigh inequality, we have
$\sigma_{\text{min}}\left(A\right)\|x\|_{2}^{2}\leq\|xA\|_{2}^{2}=xAA^{T}x^{T}\leq\sigma_{\text{max}}\left(A\right)\|x\|_{2}^{2}$
$\sigma_{\text{min}}\left(A\right)^{1/2}\|x\|_{2}\leq\|xA\|_{2}\leq\sigma_{\text{max}}\left(A\right)^{1/2}\|x\|_{2}$ (16)
Consider $A=\prod_{i=m}^{1}A_{i}$ and apply Inequality 16 recursively, we obtain
$\|x\|_{2}\prod_{i=m}^{1}\sigma_{\text{min}}\left(A_{i}\right)^{1/2}\leq\|x\prod_{i=m}^{1}A_{i}\|_{2}\leq\|x\|_{2}\prod_{i=m}^{1}\sigma_{\text{max}}\left(A\right)^{1/2}$
$\|x\|_{2}\prod_{i=m}^{1}a_{i}\leq\|x\prod_{i=m}^{1}A_{i}\|_{2}\leq\|x\|_{2}\prod_{i=m}^{1}b_{i}$
Proof of Theorem LABEL:thm:grad_bound
We first have
$\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial h_{i,t}}=1_{k}-p_{i,t},$
where $1_{k}$ is a one-hot vector over the vocabulary, the token $o_{i,t}$ has the index $k$ in the vocabulary, $p_{i,t}$ is the distribution over vocabulary with $\pi_{\theta}\left(o_{i,t}\right)=p_{i,t}(k)$.
Consider a specific layer $l$, we then have
$\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial\theta_{l}}$ $=\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial h_{i,t}}\frac{\partial h_{i,t}}{\partial a_{L}}\frac{\partial a_{L}}{\partial a_{L-1}}...\frac{\partial a_{l+1}}{\partial a_{l}}\frac{\partial a_{l}}{\partial\theta_{l}}$
$=$ $\left[1_{k}-p_{i,t}\right]W\prod_{i=l}^{L}J_{i}G_{l}.$
Applying Lemma D.1, we reach
$\|1_{k}-p_{i,t}\|_{2}a^{W}a_{l}^{G}\prod_{i=l}^{L}a_{i}^{J}\leq\|\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial\theta_{l}}\|_{2}\leq\|1_{k}-p_{i,t}\|_{2}b^{W}b_{l}^{G}\prod_{i=l}^{L}b_{i}^{J}$
We further bound
$\|1_{k}-p_{i,t}\|_{2}=\sqrt{\left(1-p_{i,t}\left(k\right)\right)^{2}+\sum_{j\neq k}p_{i,t}\left(j\right)^{2}}\geq 1-p_{i,t}\left(k\right)=1-\pi_{\theta}\left(o_{i,t}\right).$
$\|1_{k}-p_{i,t}\|_{2}$ $=\sqrt{\left(1-p_{i,t}\left(k\right)\right)^{2}+\sum_{j\neq k}p_{i,t}\left(j\right)^{2}}\leq\sqrt{\left(1-p_{i,t}\left(k\right)\right)^{2}+\left(\sum_{j\neq k}p_{i,t}\left(j\right)\right)^{2}}$
$=$ $\sqrt{\left(1-p_{i,t}\left(k\right)\right)^{2}+\left(1-p_{i,t}\left(k\right)\right)^{2}}=\sqrt{2}\left(1-p_{i,t}\left(k\right)\right)=\sqrt{2}\left(1-\pi_{\theta}\left(o_{i,t}\right)\right).$
Therefore, we further reach
$\left(1-\pi_{\theta}\left(o_{i,t}\right)\right)a^{W}a_{l}^{G}\prod_{i=l}^{L}a_{i}^{J}\leq\|\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial\theta_{l}}\|_{2}\leq\sqrt{2}\left(1-\pi_{\theta}\left(o_{i,t}\right)\right)\|1_{k}-\pi_{\theta}\left(o_{i,t}\right)\|_{2}b^{W}b_{l}^{G}\prod_{i=l}^{L}b_{i}^{J}$
Finally, by noting that
$\|\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)\|_{2}^{2}=\sum_{l=1}^{L}\|\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial\theta_{l}}\|_{2}^{2},$
hence leading to
$\frac{1}{\sqrt{L}}\sum_{l=1}^{L}\|\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial\theta_{l}}\|_{2}\leq\|\nabla_{\theta}\log\pi_{\theta}\left(o_{i,t}\right)\|_{2}\leq\sum_{l=1}^{L}\|\frac{\partial\log\pi_{\theta}\left(o_{i,t}\right)}{\partial\theta_{l}}\|_{2}$
we reach the conclusion.
Proof of Theorem 3.2
We have
$\|g_{i,t}\|_{2}=w_{it}\left|\gamma_{i,t}\right|\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}.$
Using the bound for $\|\nabla_{\theta}\log\pi_{\theta}(o_{i,t})\|_{2}$ developed in Theorem LABEL:thm:grad_bound, we reach the conclusion.
Appendix E Use of Large Language Models
In accordance with the ICLR 2026 policy, we disclose our use of Large Language Models (LLMs) during the preparation of this paper. LLMs were employed solely as writing assistants to aid in polishing the presentation, improving clarity, and shortening or rephrasing certain passages. Importantly, all scientific ideas, theoretical developments, model design, and experimental results reported in this paper are entirely conceived and executed by the authors. The LLM was never used to generate research concepts, hypotheses, or experimental findings. The authors take full responsibility for the content of the paper.